# KInIT at SemEval-2024 Task 8: Fine-tuned LLMs for Multilingual Machine-Generated Text Detection

### v1

@@ Abstract
[…]
@@ 3 System Overview
The system described in this paper. It is an ensemble using two-step majority voting for predictions, consisting of 2 LLMs (Falcon-7B and Mistral-7B) fine-tuned using the train set only, 3 zero-shot statistical methods (Entropy, Rank, Binoculars) using Falcon-7B and Falcon-7B-Instruct for calculation of the metrics, utilizing language identification and per-language threshold calibration.
PLMoE
Our initial idea representing a per-language mixture of experts. It uses Electra-Large-Discriminator for English and XML-RoBERTa-Large for each of other languages officially present in the train and dev sets. Models for languages present in the dev set only are trained using the dev set. For unknown languages the Mistral-7B fine-tuned using the whole train set is used.
rLLM2S3
The same ensemble system as LLM2S3; however, the LLMs are fine-tuned using both the train and dev sets. Also, to increase the robustness of the system, we have obfuscated 20% of the train samples during fine-tuning, by using HomoglyphAttack and inserting zero-width-joiner character, inspired by our recent work (Macko et al., 2024).
rLLM2B-ES
The post-deadline ensemble system similar to rLLM2S3; however, the Llama-2-7B is used instead of Falcon-7B and Binoculars is used solely in the statistical part (instead of a combination of 3 methods). Moreover, the fine-tuning process used the early stopping mechanism to alleviate the over-fitting.
LLM2B1
The post-deadline ensemble system using the original LLM2S3 fine-tuned Falcon-7B and Mistral-7B models; however, classification thresholds are not calibrated, but only predictions with a probability of 1 (i.e., 100% confident) are marked as machine-generated. Such predictions are combined with Binoculars zero-shot prediction using the per-language threshold calibration.

### v2

@@ Abstract
Michal Spiegel1,2and Dominik Macko1 1 Kempelen Institute of Intelligent Technologies 2 Faculty of Informatics, Masaryk University michal.spiegel@intern.kinit.sk, dominik.macko@kinit.sk
@@ 3 System Overview
PLMoE Our initial idea representing a per-language mixture of experts. It uses Electra-Large-Discriminator for English and XML-RoBERTa-Large for each of other languages officially present in the train and dev sets. Models for languages present in the dev set only are trained using the dev set. For unknown languages the Mistral-7B fine-tuned using the whole train set is used.
rLLM2S3 The same ensemble system as LLM2S3; however, the LLMs are fine-tuned using both the train and dev sets. Also, to increase the robustness of the system, we have obfuscated 20% of the train samples during fine-tuning, by using HomoglyphAttack and inserting zero-width-joiner character, inspired by our recent work (Macko et al., 2024).
rLLM2B-ES The post-deadline ensemble system similar to rLLM2S3; however, the Llama-2-7B is used instead of Falcon-7B and Binoculars is used solely in the statistical part (instead of a combination of 3 methods). Moreover, the fine-tuning process used the early stopping mechanism to alleviate the over-fitting.
LLM2B1 The post-deadline ensemble system using the original LLM2S3 fine-tuned Falcon-7B and Mistral-7B models; however, classification thresholds are not calibrated, but only predictions with a probability of 1 (i.e., 100% confident) are marked as machine-generated. Such predictions are combined with Binoculars zero-shot prediction using the per-language threshold calibration.
@@ 5 Results
5.1 Post-Deadline Study
In the post-deadline experiments (already knowing the gold labels of the test set for evaluation), we have finished Llama-2-7B model fine-tuning and retraining all three robust-version LLMs using the early stopping (to minimize the over-fitting). The results revealed that the rLlama-2 model does not suffer by over-fitting as much. Based on the test set evaluation and by examining various combinations, the retrained rLlama-2-ES and rMistral-ES seemed like good candidates to combine with Binoculars (rLLM2B-ES), outperforming the winning system in the competition.
Early stopping helped a lot in boosting performance generalization (i.e., reducing over-fitting) of our per-language mixture-of-experts ensemble system (PLMoE-ES), achieving one of the highest AUC ROC using the test set. Nevertheless, in the accuracy as an official metric, it would not outperform the other system alternatives.
In addition, we have noticed that optimal thresholds for fine-tuned LLMs are often set to 1.0 by using purely the dev set samples machine-class probabilities. Therefore, we have fixed the thresholds to 1.0 for the LLM2B1 system (containing only models we have trained before the deadline), meaning that the machine-class predictions of the LLMs are used only when having 100% confidence (otherwise considered human-written). Such predictions, when combined with Binoculars, achieved even higher performance using the test set data (0.9708). Thus, we had such a system trained before the deadline; however, we have not noticed such a threshold bringing the best performance in time. Moreover, when looking at the accuracy for the dev set, we do not see why we would select such a system for the submission. It can be just a coincidence that it performs so well using the test set data. Further experiments are required to examine this phenomenon using independent out-of-distribution data.
Also, even when our rLLM2B-ES system alternative or the rMistral1 single-model system would won the competition, we are now not sure that we would be confident enough (about not being over-fitted) to submit it as the final system without evaluation on the external dataset. Thus, we have submitted best what we could at the time.
5.2 Per-language Analysis
For a deeper insight of the proposed system (LLM2S3) performance, we have performed an analysis per each language identified in the test set. The results are provided in Figure 2. Interesting is that it achieved the highest accuracy for the Italian surprise language (it). Lower accuracy is evident for German and Arabic languages, although present in the dev set. It must be noted that this version of the system was not trained using the dev set, only the classification threshold calibration used such data. Therefore, the robust versions of system alternatives are expected to provide higher performance especially in those languages.
Figure 2: Per-language test-set performance (it is a surprise language, de and ar are in the dev set only). Axis scale for Accuracy is shown from 0.9 to 1.0. The per-class samples counts are provided in the top-right table.
6 Conclusion
To cope with the problem of multilingual, multidomain, and multigenerator machine-generated text detection, we have proposed an ensemble system using 2 LLMs (Falcon-7B and Mistral-7B) fine-tuned for the binary sequence classification task. We have further combined the predictions with the statistical metrics of Entropy, Rank, and Binoculars using a two-stage majority voting. The classification thresholds in our system have been calibrated in a per-language manner, for which we have utilized the FastText language identification. A combination of fine-tuned LLMs and statistical detection seems to be the right way to cope with generalization of the detection performance. Out of the evaluated single-model systems, Mistral-7B is the best candidate for fine-tuning, which by itself can bring a remarkable classification performance. Further improvements of the system could be easily achievable by hyper-parameters optimization, which we have not done in the submitted system due to lack of time.
Acknowledgements
This work was partially supported by the projects funded by the European Union under the Horizon Europe: AI-CODE, a project funded by the European Union under the Horizon Europe, GA No. 101135437, and VIGILANT, GA No. 101073921. Part of the research results was obtained using the computational resources procured in the national project National competence centre for high performance computing (project code: 311070AKF2) funded by European Regional Development Fund, EU Structural Funds Informatization of Society, Operational Program Integrated Infrastructure.
Appendix A Computational Resources
For experiments regarding model fine-tuning and inference processes, we have used $1\times$1 × NVIDIA GeForce RTX 3090 24GB GPU and $1\times$1 × A100 40GB GPU, cumulatively consuming around 10,000 GPU-core hours. For combining the results and analysis, we have used Jupyter Lab running on 4 CPU cores, without the GPU acceleration.